Use the Laplace transform to solve the initial-value problem y - 8 y + 20y = te, y(0) = 0, y(0) = 0
Answer #1

Sol +e} = 20 + أو 8-او 0=والا ,0 (0 الا (ی ن = ای و 12 (5) - (sys = والا 12 رهان - (هالا 5- (sys = ( و 12 :: { 4 *** * (4) 90 =(1-8) A + B +D+2C. o= -9AAB te. 02-91-C) + 3 + D.-20. 10 = 70+07 29 PE so 1 A=.6 169/ B- Ķ, C- 169 169. -65+29 2 1 6 16915